OpenStudy (anonymous):

find the equation of the a line passing through the point (2, 1, 3) and perpendicular to the line (x-1)/1=(y-2)/2=(z-3)/3 and x/-3=y/2=z/5

OpenStudy (amistre64):

find the equation of the a line passing through the point (2, 1, 3) and perpendicular to the line (x-1) (y-2) (z-3) --- = ---- = ---- 1 2 3 and x y z --- = --- = --- -3 2 5

OpenStudy (anonymous):

need the direction ratios of the line using the condition of perpendicularity, l+2m+3n=0 -3l+2m+5n=0

OpenStudy (amistre64):

cross your vector parts to get the normal vector

OpenStudy (anonymous):

(l,m,n) are the direction ratios of the required line

OpenStudy (amistre64):

<1,2,3> x <-3,2,5> = normal vector right?

OpenStudy (amistre64):

n<a,b,c> ; at point P(x,y,z) x-Px y-Py z-Pz ----- = ----- = ----- gets your line a b c

OpenStudy (anonymous):

(l,m,n)=(2,-7,4)

OpenStudy (anonymous):

line is x-2/2=y-1/-7=z-3/4

OpenStudy (amistre64):

i can make a matrix in my ti83, but cant figure out how to math it in lol

OpenStudy (amistre64):

\[\left|\begin{array}c i&j&k\\1&2&3\\ -3&2&5 \end{array}\right|\] \[(2.5-2.3)i - (1.5-(-3).3)j+(1.2-(-3.2))k\] \[4i-14j+8k\] \[2<2,-7,4>\] right?

OpenStudy (amistre64):

yay!! ... we match

OpenStudy (anonymous):

:P

OpenStudy (amistre64):

another way to write the line equation is: x = 2 +2t y = 1 -7t z = 3 +4t

OpenStudy (anonymous):

yes, symmetric form :)

OpenStudy (anonymous):

thank u very much